OpenStudy (anonymous):

1. What is the quotient x^2-16 over 2x^2-9x+4 all divided by 2x^2+14x+24 over 4x+4? A.1 over x+3 B.2x+2 over x+3 C.2x+2 over 2x^2+5x-3 D.2(x+1) over 2x^2-5x-3 Medal for the best answer.

OpenStudy (anonymous):

((x-4)(x+4)/(2x-1)(x-4)) / (2(x+3)(x+4))/(4(x+1)) (x+4)/(2x-1) / 2(x+3)(x+4) / (4(x+1)) (x+4)/ (2x-1) x (4(x+1)/ (2(x+4)(x+3) R: 4/2 = 2 and (a/b)/(c/d) = (a/b)x(d/c) (x+4) / ( 2x-1) x 2(x+1) / (x+4) (x+3) R: (x+4)/(x+4) = 1 Answer: 2(x+1) / (x+3)(2x-1) = 2(x+1) / 2x^2 + 5x -3